Use inductive reasoning to predict the next complete line in the pattern.
55 - 6 = 49
565 - 67 = 498
5665 - 678 = 4987
Please help me. I have no idea where to even being

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(57290)

predict the next complete line in the pattern.
55 - 6 = 49
565 - 67 = 498
5665 - 678 = 4987
============================
56663 - 6789 = 49876
=============================
Cheers,
Stan H.

Answer by solver91311(16877)


First number, first calculation, zero 6s.
First number, second calculation, one 6
First number, third calculation, two 6s.
First number, fourth calculation, hmmm, what do you think?

Second number, first calculation, 6
Second number, second calculation, 67
Second number, third calculation, 678
So what do you think for this one?

Last pattern
49
498
4987
4987? What do you think replaces the (?)?

John